A goal is scored in hockey when the puck, a vulcanized rubber disk 2.5 cm (1 in.) thick and 7.6 cm (3 in.) in diameter, is driven into the opponent's goal. Find the volume of a puck in cubic centimeters and cubic inches.
Answer by cleomenius(959)

We will do inches first.
Volume = area * height.
area = pi*r^2
r = 1.5, Area = 1.5^2 *pi Area = 2.25*pi
Volume = 2.25pi inches^2 * 1 in = 2.25pi cubic inches.
Now cm.
r = 3.8 cm area = 3.8^2 * pi = 14.44*pi
Volume = 14.44pi cm^2 * 2.5 cm = 36.1*pi cm cubed.
Cleomenius.